A woman of mass m = 60 kg descends in an elevator in that briefly accelerates downward at a = 0.20g. If she stands on a scale, what is the reading on the scale that is apparent/effective weight w?


1
Expert's answer
2021-03-18T15:34:28-0400

Solution:


When the elevator is not moving, the scale reads 60*9.8 = 588N


If the elevator is falling acceleration 9.8, the scale reads zero


W = m*(g+a) = 60(9.8 - 0.2*9.8) = 470N